# FormGym: Doing Paperwork with Agents

### v1

@@ 3.4 Evaluation
Each field is also associated with a correctness function to provide fair evaluation of answers with multiple correct formats, such as telephone numbers. If a field contains multiple text inputs, we concatenate them. We choose field accuracy as our primary evaluation metric, ignoring those that should be empty according to the ground truth label to avoid inflating accuracy. A text input is considered to be inside a field if the center point of the text is within a designated bounding box.
3.5 Baseline Agents
We experiment with both classic VLAs and GUI agents capable of interacting with browser and desktop applications.
3.5.1 Vision Language Models
We prompt VLAs with API documentation, examples of all available actions, and a natural language descriptions of the user profile (Appendix A.4).
3.5.2 GUI Agents
We instantiate GUI agents Claude Computer Use and OpenAI Operator with the free in-browser photo editing application Photopea111photopea.com, whose interface is nearly identical to Photoshop (Appendix A.3). We prompt GUI agents with natural language user profile descriptions and instructions to complete the form. For accessibility and cost reasons, we limit operators to five minutes per page. Prompts include detailed instructions on how to use the Photopea interface, without which GUI agents fail completely (Appendix A.5).
4 FieldFinder
We observe that large baseline VLAs make coherent API calls, but universally struggle to place text in appropriate locations. To ameliorate this issue, we create the FieldFinder tool. FieldFinder takes a form image and text description of the name of the target field as input and predicts the bounding box around the valid input space (Figure 2).
Figure 2: Agent use of the FieldFinder tool. 1) The agent ingests an input form or database. 2) The agent requests the location of an empty field by name. 3) The FieldFinder returns the bounding box around the target field to the agent.
4.1 Dataset
To train the FieldFinder tool, we create a (Document, target field name, bounding box) dataset using question/answer relations in the FUNSD and multilingual XFUND Xu et al. (2022) form understanding datasets. Since FUNSD and XFUND forms contain responses in answer fields, we use horizontal inward content aware fill 222github.com/light-and-ray/resynthesizer-python-lib to automatically remove text while generally preserving formatting such as lines and table boundaries.
Table 1: Generated by Spread-LaTeX
Forms NL Fields DB Fields Users
Auto Loans 4 357 0 4
Database 2 10 39 1
FUNSD 50 50 0 1
Table 2: Total form pages, fields whose values are supplied in natural language, supplied in a dataabase, and user profiles in FormGym tasks.
4.2 Training
We fine-tune a Florence 2 Large Xiao et al. (2024) vision foundation model to predict the answer bounding box coordinates given the target question text and document. We choose Florence 2 because its pretraining contains both open-vocabulary object detection and tasks requiring OCR, minimizing the distribution shift between pretraining and fine-tuning. Florence 2 Large has only 0.77B parameters, contributing minimal latency and memory overhead when augmenting with much larger VLAs. We train the FieldFinder for 4 epochs using early stopping, batch size 8, learning rate 1e-6 on 1x NVIDIA A100 GPU for approximately 20 hours. The FieldFinder achieves an intersect-over-union of 20.9% on the FUNSD test set.
5 Results
Auto Loans (Text) Auto Loans (Doc Transfer) Database FUNSD
One-shot Iterative One-shot Iterative Iterative One-shot
Aria 25B 0.0 0.1 0.0 0.0 1.0 0.0
Claude 3.7 0.1 0.3 0.2 0.2 2.7 32.0
GPT-4o 0.6 0.6 0.0 0.6 0.0 2.0
Llava 7B 0.0 0.0 0.0 0.0 0.0 0.0
Molmo 7B 0.0 0.0 - - 0.0 0.0
Aria 25B + FL (ours) 6.2 6.7 1.4 2.4 1.0 28.0
Claude 3.7 + FL (ours) 18.8 14.9 5.8 7.2 32.0 52.0
GPT-4o + FL (ours) 12.2 17.2 7.1 6.4 0.0 56.0
Llava 7B + FL (ours) 1.5 0.4 0.4 0.0 1.0 6.0
Molmo 7B + FL (ours) 0.4 0.0 - - 20.0 20.0
OpenAI Operator - 18.3 - 3.9 36.0 50.0
Claude Computer Use - 10.6 - 1.4 44.0 68.0
Table 3: Average form completion percentage (correct fields / all fields). Iterative FUNSD is omitted because FUNSD forms contain only one empty field. One-shot Database is omitted because at least two turns are necessary. Molmo is not trained for multi-image prompting
Overall, VLAs struggle with this task, with models performing best on FUNSD and worst on Database (Table 3). Baseline models generally score $\leq$ 1%, except for Claude on FUNSD and Database (32% and 2.7%, respectively). When introducing FieldFinder, we observe equal or better performance in all cases. In the best case, GPT-4o’s performance on FUNSD increases from 2% to 56%. We observe smaller gains, up to 16.9 percentage points on Auto Loans (GPT-4o), and 29.3 points on Database (Claude 3.7). Certain small, open-source models including Aria 25B and Molmo 7B achieve significant performance improvements with FieldLocalizer. GPT-4o and Claude also struggle to chain actions in the more complex Doc Transfer and Database tasks. GPT-4o performs especially poorly, suggesting the user query the database herself, then signing a page footer with "Your Name".
Across all tests, GUI agents performed as comparable or better than VLAs, except in Doc Transfer. Although GUI agents still made localization errors, these were typically less distant than those of VLAs. GUI agents often did not complete the Auto Loans and Database tasks within the 5 minute timeframe, negatively impacting completion. Although Claude Computer Use was more accurate than OpenAI Operator, it performed actions about half as fast, bottlenecking completion.
6 Discussion
We attribute weak baseline model performance to several failure modes. The inability to localize answer fields and chain actions are the primary weaknesses in Claude and GPT-4o. Although Auto Loans contains 357 graded fields, Claude and GPT-4o make as few as 71 placement attempts in some cases, suggesting a failure in document understanding and completeness tracking. Claude and GPT-4o also struggle to recover from mistakes. Although they are provided with an API to delete text, its usage is vanishingly rare.
When using FieldFinder, accuracy on FUNSD is uniformly higher than on other tasks. We attribute the performance discrepancy to several factors. First, FieldFinder was trained on FUNSD, so testing on Database and Auto Loans represents a significant distribution shift in inputs. Second, Auto Loans requires differentiating between relationally complex fields, such as "applicant first reference name" versus "co-applicant second reference name", indicated by physically distant table headers. To model the upper limit of the impact of FieldFinder, we conduct an ablation study wherein models are prompted with the exact centroid coordinates of fields. Under these conditions, GPT-4o achieves 77% accuracy and Claude 3.7 achieves 82%, suggesting field localization errors account for about 4/5 errors, while document understanding accounts for the other 1/5. Future work should explore training field localizers on a broader distribution of documents and improving foundational models’ visual reasoning and backtracking abilities.
Given GUI agents accurate but sluggish performance, future research should prioritize inference speed and UI generalization with a minor focus on localization. Poor inference efficiency also raises costs, which we calculate to be approximately \$1 USD per Auto Loans page. We note that without iterative and specific prompt engineering, GUI models perform no successful actions.
@@ 7 Conclusion
We present a challenging agent benchmark for image-domain form filling and contribute a cross-model field localization tool that can retrofit VLAs, increasing form completion by up to 54 percentage points with minimal overhead.
@@ 8 Limitations
For cost and accessibility reasons, this benchmark only assesses performance on a small sample of commercial, English, single-page documents in the image domain. PDF features, such as attachments, page manipulation, passwords, interactive fields, and editing are also not evaluated.
Because text placement accuracy is determined by whether its geometric center is contained within a field, the text itself may sometimes overflow the field boundary and still be marked as correct. Although aesthetically unpleasing, we observe that these placements would generally be comprehensible to human readers.
@@ 9 Ethical Considerations
The validity and legal status of electronically or agent-generated signatures is complex and varies between jurisdictions. We recommend that automated signature placement only be used as a suggestion rather than a fully automated process. Similarly, due to the legal weight of many forms, we recommend that all agent-filled forms be proofread by a qualified human prior to submission.
Appendix A Appendix
A.1 Example Output
{adjustwidth}
-2cm-2cm
Figure 3: Output by Claude 3.7 in the Auto Loans One-shot task. Baseline (top), with FormFiller (middle), with ground truth field centroids in prompt (bottom).
A.2 Additional Experimental Details
In iterative flow, after each agent turn, the agent is presented with an updated document reflecting any text or signatures it placed. On subsequent turns, we provide the agent with the following feedback for each action in the prompt:
PlaceText
- Whether the text was placed successfully and where
DeleteText
- What text was deleted, if any
SignOrInitial
- Whether the signature was placed successfully and where
QuerySql
- The SQL output or error message
Intersection over Union (IoU) is calculated as:
IoU=Area of OverlapArea of $\text{IoU}=\frac{\text{Area of Overlap}}{\text{Area of Union}}=\frac{|A\cap B|}{|A\cup B|}$
Where:
$A$ = predicted bounding box
$B$ = ground truth bounding box
$|A\cap B|$ = area of intersection
$|A\cup B|$ = area of union
In our FUNSD dataset, the ground truth bounding box is taken to be the envelope of the answer string, which is generally a subset of the actual field. This may contribute to an underestimate of actual IoU accuracy and FUNSD placement accuracy. However, because we care predominately about the centroid of the placement text, training and predicting a smaller bounding box contained within the actual field should not negatively impact training.
A.3 GUI Agent Implementation Details
We the performance between Anthropic Claude Computer Use, and OpenAI Operator. To equip models with the necessary tools, we set up an environment to allow models to place text on the a PNG version of the form through the online graphic editor Photopea. This enables agents (notably OpenAI Operator) to operate entirely in a web browser environments. Instructions on how to use the system were provided specifically to isolate out the form filling performance and remove confounding factors with interface-use performance. We also gave specific interface instructions to prevent models from leaving the tab or deleting the form.
In the Doc Transfer task, a reference document was loaded in another tab inside Photopea. Instructions in the prompt were adjusted to account for the reference document.
We provided GUI agents a REPL connect to the database in Google Colab with ipywidgets for in-browser querying.
A.4 Example VLA Prompt
@@
The user’s reference’s cell phone is: 415-555-$1$
The user’s reference’s home phone is: 415-555-$5$
@@
The joint filer’s reference’s cell phone is: 414-555-$9$
@@
The user’s second reference’s cell phone is: 241-444-$4$
The user’s second reference’s home phone is: 241-222-$2$
@@ 3. Proceed to step 1 as you will remain in text edit mode
A.6 Additional Results
Auto Loans (Text) Auto Loans (Doc Transfer) Database FUNSD
One-shot Iterative One-shot Iterative Iterative One-shot
Aria 25B 0.0 0.2 0.0 0.0 0.1 0.0
Claude 3.7 0.2 0.3 0.4 0.7 5.6 34.0
GPT-4o 0.7 0.9 0.0 1.3 0.0 0.8
Llava 7B 0.0 0.0 0.0 0.0 0.0 0.0
Molmo 7B 0.0 0.0 - - 0.0 0.0
Aria 25B + FL (ours) 14.6 5.4 4.0 2.3 0.3 3.2
Claude 3.7 + FL (ours) 23.8 19.3 18.9 27.0 47.8 51.0
GPT-4o + FL (ours) 21.6 22.1 23.2 15.7 0.0 30.4
Llava 7B + FL (ours) 19.3 5.8 2.9 0.0 4.8 1.3
Molmo 7B + FL (ours) 3.9 0.0 - - 10.4 1.8
OpenAI Operator - 59.4 - - 81.8 50.0
Claude Computer Use - 86.8 - - 100.0 68.0
Table 4: Placement accuracy (correct placements / total placements)
Auto Loans (Text) Auto Loans (Doc Transfer) Database FUNSD
One-shot Iterative One-shot Iterative Iterative One-shot
Aria 25B 0.0 0.5 0.0 0 0.3 0.0
Claude 3.7 0.5 1.0 0.8 0.75 0.7 16.0
GPT-4o 2.0 2.0 0.0 2 0.0 1.0
Llava 7B 0.0 0.0 0.0 0 0.0 0.0
Molmo 7B 0.0 0.0 - - 0.0 0.0
Aria 25B + FL (ours) 22.0 23.8 5 8.5 0.3 14.0
Claude 3.7 + FL (ours) 67.0 53.3 20.75 25.75 8.0 26.0
GPT-4o + FL (ours) 43.5 61.3 25.3 23 0.0 28.0
Llava 7B + FL (ours) 5.5 1.3 1.3 0 0.3 3.0
Molmo 7B + FL (ours) 1.3 0.0 - - 5.0 10.0
OpenAI Operator 65.3 14 9 25
Claude Computer Use 37.75 5 11 34
Table 5: Average Total Correct Fields
Auto Loans (Text) Auto Loans (Doc Transfer) Database FUNSD
One-shot Iterative One-shot Iterative Iterative One-shot
Aria 25B 150.0 282.5 85.8 205.25 177.5 649.0
Claude 3.7 300.0 303.3 166.8 102.25 12.0 47.0
GPT-4o 274.0 228.0 71.0 156 3.0 127.0
Llava 7B 27.8 13.0 14.8 22.5 7.8 217.0
Molmo 7B 6.5 31.0 - - 75.3 472.0
Aria 25B + FL (ours) 151.0 440.8 125.5 365 98.5 443.0
Claude 3.7 + FL (ours) 281.3 276.3 109.5 95.5 16.8 51.0
GPT-4o + FL (ours) 201.5 277.0 109.0 146.75 3.0 92.0
Llava 7B + FL (ours) 28.5 23.0 43.0 12.25 5.3 234.0
Molmo 7B + FL (ours) 32.0 2.0 - - 48.0 552.0
OpenAI Operator - 110 - 25.5 11 50
Claude Computer Use - 43.5 - 16 11 50
Table 6: Average Total Incorrect Placements

### v3

@@ 6.2 Metrics
We evaluate field completion correctness based on whether each field contains the correct value, as defined in Sections 3.1 and 3.2. If a field contains multiple text inputs, we concatenate them with a space. We choose field accuracy as our primary evaluation metric, ignoring those that should be empty according to the ground truth label to avoid inflating accuracy. A text input is considered to be inside a field if the center point of the text bounding box is within the field. A side effect of this rule is the increased difficulty of forms with numerous fields. On such forms, inaccurate placements are more likely to fall into unrelated fields, resulting in a double error.
We choose center point correctness instead of other metrics including intersect-over-union or full enclosure because it does not rely on a priori knowledge of text size for models to accurately place text. It also avoids ambiguity in how to judge text that slightly overhangs out of a field or into another field when, to a human, the intention of the form filler would be obvious.
6.3 Performance Comparison
Overall, VLAs struggle on FormGym, with models performing best on Form-NLU and worst on Auto Loans (Table 2). Baseline VLAs generally score $\leq$ 3%. Set of Marks VLA augmentation offers no improvement over baseline VLA setup. Claude Comptuer Use, however, achieves an average accuracy of 21.1% Our method using Claude 4 + FieldFinder, achieves the highest accuracy at 23.0%.
One outlier is baseline VLA Claude 4’s performance on FUNSD, achieving 21.0%. It appears that, in some cases, Claude can performing vertical but not horizontal localization, frequently generating $x=0.5$. Because FUNSD fields are very wide, this results in higher accuracy. We observe a similar effect in Claude Computer Use, which achieves 60.2% accuracy. However, when introducing FieldFinder, we observe equal or better performance in all cases (Figure 2).
(a) Baseline
(b) With FieldFinder
(c) Ground Truth
Figure 5: Example output by Claude 4 baseline, with FieldFinder (ours), and ground truth in the Auto Loans (Text) One-Shot task. We attribute FieldFinder’s leftward bias to supervision artifacts: training labels mark left-biased value text rather than full fields. Without FieldFinder, Claude appears to struggle more with horizontal spacing that with vertical spacing, assigning most placements an x coordinate of exactly 0.5 (not centered due to figure cropping).
Augmenting with FieldFinder, Claude 4 improves from 4.5 to 23.0% average accuracy. In the best case, Claude 4’s performance on Form-NLU increases from 0% to 54%. We observe smaller gains, up to 9.2% percentage points on Auto Loans (Text) (Figure 5). Notably, Claude 4 with FieldFinder surpasses Claude Computer Use on both digital-born datasets, Auto Loans and Form-NLU. All open-source models also achieve performance improvements with FieldFinder, with Aria leveraging the tool better than smaller models. GPT-5 and Claude appear to struggle to chained reasoning in the more complex Auto Loans (Image) task, where accuracy falls by around 50% compared to when source information is supplied in text form.
We illustrate the most common FieldFinder errors in Figure 6. Primary failure modes include placing text slightly outside the field and placing text in the wrong field with a lexically similar name. The later case can generate secondary failure by adding additional text into the incorrect field, sometimes superimposing it on the otherwise correct answer.
Our FieldFinder localization tool demonstrates a method for equipping zero-shot VLAs with field localization abilities. Without FieldFinder, VLAs struggle to even approach this task due to a lack of training data involving pixel-level predictions, with frontier models scoring below 1% accuracy in multiple domains. As expected, FieldFinder performs strongest on high quality English document images with relatively few fields. FieldFinder uses only 0.77B parameters, requiring negligible memory and latency overhead when augmenting both closed- and open-source models with no need for additional training.
We note that XFUND accuracy falls substantially below the English dataset trend line, despite constituting the largest portion of training data. We attribute this to the multi-lingual nature of the XFUND dataset, a well studied weakness in models pre-trained primarily on English data Geigle et al. (2025).
7 Error Analysis
Surprisingly, we see little improvement between One-Shot and Iterative flows, except in the case of Aria, which shows a 1.27-fold increase in accuracy (Table 2). Analyzing VLA trajectories, we attribute this to frontier models calling the Terminate action after the first turn in most (or in the case of Claude, all) trajectories. Although these models are prompted with the number of remaining opportunities to edit the form, they appear to lack the self-awareness to doubt their own accuracy and utilize future turns for error correction.
Figure 6: Example failure modes using FieldFinder in an example from Form-NLU. The answer for Field 1 was inaccurately superimposed on Field 3 which is otherwise correct. Answers for Fields 2 and 4 were transposed. Answers in fields 5-7 used an incorrect date format. Field 8 is correct. The answer for Field 9 is outside the field. The answer for Field 10 is placed in Field 11.
GUI agents present one path to end-to-end form filling, with Claude Computer Use achieving 21.1% accuracy. Qualitatively, we observe that it is particularly challenged by long documents, often failing to finish within the allotted time. It also typically fails on forms containing tables, resulting in only 2.7% accuracy on Auto Loans (Text). Overall, Claude Computer Use falls short of our Claude 4 + FieldFinder method (23.0%). Moreover, Claude Computer Use costs \$5.40 per 1000 fields, compared to \$0.043 for Claude 4 + FF. Furthermore, Claude Computer Use requires 28 minutes per 100 fields, compared to approximately 20 seconds seconds for Claude 4. Ultimately, we find that Claude Computer Use requires over 100x the cost and 84x the time to complete forms while providing worse overall accuracy.
Figure 7: Error rates from generation (VLA performance given ground truth coordinates), document understanding (difference between number of placements and fields), and localization (FieldFinder accuracy given ground truth field names). Localization error shown applies only to our FieldFinder pipeline, whereas generation and document understanding errors apply to both FieldFinder and baseline VLA pipelines instantiated with Claude 4.
We conduct studes to attribute errors to either the document understanding or document reasoning steps of the VLA pipeline using Claude 4. Document understanding errors arise from VLAs failing to understand the input document image. Document reasoning errors, in contrast, arise from erroneous generations by the VLA even when provided with ground truth free text descriptions of field names and locations. Compared to the localization errors discussed in Section 4.3, these errors are somewhat less common (Figure 7) overall, but more common in the high quality, low field density Form-NLU dataset.
7.1 Document Reasoning Error
To study the influence of VLA document reasoning errors, we provide the overall strongest model, Claude 4, with ground truth coordinates to the centroid of each field and its name. Under these conditions, the VLA no longer needs to localize suggested fields, must only map information in the context to the appropriate bounding box. In general, accuracy is relatively good, ranging from 74% Auto Loans (Text) to 83% (FUNSD). However, accuracy drops from 74% to 49% when the source persona is provided in image form (Figure 7). This indicates that, although Claude can understand documents well in simple tasks, it struggles to apply these abilities when they are integrated into a more complex workflow.
7.2 Document Understanding Error
In a third failure mode, models may fail to detect fields entirely, resulting in neither a localization nor a reasoning error. In Figure 7, we plot document understanding errors as the absolute difference between the number of text placements and the number of fields. We observe that Claude and GPT-5 generate as few as 0.42 placements per field in the Auto Loans task, indicating they did not even attempt to fill more than half of fields. This tendency is compounded by a frequent early termination rate in Iterative tasks, resulting in a relatively modest improvement between One-Shot and Iterative tasks. Furthermore, despite being equipped with the delete action, we find that its use is vanishingly rare.
7.3 Error Comparison
We observe that even with FieldFinder, localization remains the largest source of errors in all datasets except for Form-NLU. We note that these failure modes are not well-correlated, suggesting that future work may require a multi-faceted approach. Additionally, language does not appear to affect document understanding or reasoning significantly, with comparable results across XFUND and FUNSD. Therefore, we encourage future work addressing document understanding, reasoning, and localization in VLAs.
@@ 8 Conclusion
We present a large and realistic agent benchmark for systematically evaluating end-to-end form filling across multiple languages and domains. We also contribute the FieldFinder tool that enables VLAs to overcome their main bottleneck in field localization. With FieldFinder, frontier models improve performance across all domains, with an average increase from 5.6% to 27.0% accuracy.
@@ 9 Limitations
This benchmark only assesses single-page documents in the image domain. PDF features, such as attachments, page manipulation, passwords, interactive fields, and editing are also not evaluated.
We omit validation on other GUI agents (e.g. OpenAI Operator) due to insufficient quota offerings at time of publication.
Because text placement accuracy is determined by whether its geometric center is contained within a field, the text itself may sometimes overflow the field boundary and still be marked as correct. Although aesthetically unpleasing, we observe that these placements would generally be comprehensible to human readers. On FUNSD, XFUND, and Form-NLU, accuracy is defined by the bounding box of keys themselves, as opposed to the field containing them. This deflates accuracy in some cases, motivating our inclusion of the Auto Loans dataset.
@@ 10 Ethical Considerations
The validity and legal status of electronically or agent-generated signatures is complex and varies between jurisdictions.
@@
We recommend that automated signature placement only be used as a suggestion rather than a fully automated process. Similarly, due to the legal weight of many forms, we recommend that all agent-filled forms be proofread by a qualified human prior to submission.
Appendix A Annotation Scheme
We annotate all bounding boxes requiring applicant input, ignoring those reserved for loan officers. Bounding boxes extend to the border of the suggested text input area. For table cells, this corresponds to the entire cell. For underlines, this corresponds to the underlined area. For unmarked suggested fields (e.g. colon), we set the height equal to the suggesting text and the width equal to the area between the field name and the next element to the right. We ensure fields never overlap. Observing that all fields in the Auto Loans dataset are rectangular, we use only orthogonal rectangular bounding boxes. We label each bounding box with its desired information (e.g. Full Name). For each bounding box label, we write a correctness function that returns true or false as a function of the input text and the source persona. In general, we attempt to accept as many reasonable variations in style, punctuation, formatting, and capitalization as possible.
Appendix B Form-NLU Synthesis Process
Due to an apparent corruption in the Form-NLU dataset, we were not able to extract relations. Instead, we recacluated relations using GPT-5 as an OCR model. We prompted GPT-5 with the document, instructing it to produce all key-value pairs, including hierarchical information. We then rejected any key value pairs that did not appear exactly in the document. Finally, using original Form-NLU data, we include annotations containing keys, values, and field boundaries.
Appendix C Comparison of Existing Methods
We compare existing methods in Table 3. Prior work falls primarily into four categories: VLAs and VLA augmentations including Set of Marks; GUI agents; Document segmentation models; and commercial tools.
Method Image-Only Paper Forms Empty Fields Generative Open Source
VLA
VLA + Set of Marks
WebVoyager
OmniParser
PubLayNet $\sim$
DocLayNet $\sim$
LayoutLM $\sim$
SelfDoc $\sim$
DocFormer $\sim$
FormNet $\sim$
Mac OS Preview $\sim$
Amazon Textract $\sim$
GUI Agents $\sim$ $\sim$
VLA + FieldFinder (Ours)
Table 3: Comparison of existing methods. Definitions: Generative - The method can do open-domain text generation and follow instructions. Empty Fields - The method has a framework to handle empty fields, including those defined by underlines, table cells, colons, and check boxes, and blank space. Image-Only - The method can operate on images without the need for structured or metadata, such as a DOM or rich text PDF. Open Source - The method is open source. Paper Forms - The method is pretrained on images of paper forms, rather than GUIs. GUI agents includes Shen et al. (2024a); Qin et al. (2025a); Liu et al. (2024); Hong et al. (2024); Wei et al. (2025); Qi et al. (2025); Ma et al. (2023); Putta et al. (2024); Shen et al. (2024b) ✓: satisfies. $\sim$: framework handles some but not all types of empty fields.
Appendix D Bounding Box Evaluation
In the case of Form-NLU, XFUND, and FUNSD examples, defining text boxes by the word bounding boxes may underestimate the size of the input field when words do not entirely cover the field. However, this underestimation is roughly compensated by defining placed text by its centroid, rather than its bounding box. Under this definition, text overflowing from the bounding box is not penalized, as long as the cetroid is inside the bounding box. Bounding boxes are defined by the field itself in auto loans forms, meaning only the latter bias exists. However, in such cases, it is guaranteed that at least 1/4th of text placements marked as accurate falls within the text box. Qualitatively, we observe that, although aesthetically unpleasing, it is generally clear which field these values apply to. We note however, a bias in GUI agents, which tend to place text in the top-right of English forms, which tend towards a leftward bias in FUNSD and Form-NLU. We therefore evaluate GUI agent generations manually, in order to provide a more faithful representation of their accuracy. We find minimal difference between manual and automatic evaluations on VLA generations, so report the automatic generations because they are more replicable.
Appendix E Example VLA Prompt
@@
The user’s reference’s cell phone is: 415-555-1111
The user’s reference’s home phone is: 415-555-5555
@@
The joint filer’s reference’s cell phone is: 414-555-9999
@@
The user’s second reference’s cell phone is: 241-444-4444
The user’s second reference’s home phone is: 241-222-2222